Lab 1 Schedule

• The lab will be in Room 216 of the old engineering building

1. Monday (2/3/03) • 3:00-3:20: Hewitt, Hill, Ritchson

• 3:20-3:40: Bush, Morgan Nwosisi

• 3:40-4:00: Ferguson, King, Leavesley • 4:00-4:20: Cockrell, Joseph, Scheffler • 4:20-4:40: Andrews, Clark, Hawkins

2. Wednesday (2/5/03) • 3:00-3:20: Bolden, Calhoun, Dupree • 3:20-3:40: Dawson, Ferguson, Lala • 3:40-4:00: Bullion, Fritz, June

• 4:00-4:20: Cambrice, Camp, Nadaskay • 4:20-4:40: Mertz, Jean, Johnson